if i have a triangle with the angles A=40 B=50 and C=90 and side c=5 what are a and b

Answer by stanbon(75887)

if i have a triangle with the angles A=40 B=50 and C=90 and side c=5 what are a and b
----
sin(B) = b/5
b = 5*sin(50) = 3.83..
---
a = 5*sin(40) = 3.21
============================
Cheers,
Stan H.
================